how find ph values

Grade:9

1 Answers

Anurag Kishore
13 years ago

Hi

pH

pOH

pH is a measure of the hydrogen ion concentration, [H+]

pOH is a measure of the hydroxide ion concentration, [OH-]

pH is calculated using the following formula:
pH = -log10[H+]

pOH is calculated using the following formula:
pOH = -log10[OH-]

Example 1:

Find the pH of a 0.2mol L-1
(0.2M) solution of HCl

  • Write the balanced equation for the dissociation of the acid
    HCl -----> H+(aq) + Cl-(aq)
  • Use the equation to find the [H+]:
    0.2 mol L- HCl produces 0.2 mol L-1 H+ since HCl is a strong acid that fully dissociates
  • Calculate pH: pH = -log10[H+]
    pH = -log10[0.2] = 0.7

Example 1:

Find the pOH of a 0.1mol L-
(0.1M) solution of NaOH

  • Write the balanced equation for the dissociation of the alkali
    NaOH -----> OH-(aq) + Na+(aq)
  • Use the equation to find the [OH-]:
    0.1 mol L-1 NaOH produces 0.1 mol L-1 OH- since NaOH is a strong alkali that fully dissociates
  • Calculate pOH: pOH = -log10[OH-]
    pOH = -log10[0.1] = 1

Example 2:

Find the pH of a 0.2 mol L-1
(0.2M) solution of H2SO4

  • Write the balanced equation for the dissociation of the acid
    H2SO4 -----> 2H+(aq) + SO42-(aq)
  • Use the equation to find the [H+]:
    0.2 mol L-1 H2SO4 produces 2 x 0.2 = 0.4 mol L-1 H+ since H2SO4 is a strong acid that fully dissociates
  • Calculate pH: pH = -log10[H+]
    pH = -log10[0.4] = 0.4

Example 2:

Find the pOH of a 0.1mol L-1
(0.1M) solution of Ba(OH)2

  • Write the balanced equation for the dissociation of the alkali:
    Ba(OH)2 -----> 2OH-(aq) + Ba2+(aq)
  • Use the equation to find the [OH-]:
    0.1mol L-1 Ba(OH)2 produces 2 x 0.1 = 0.2 mol L-1 OH- since Ba(OH)2 is a strong alkali that fully dissociates
  • Calculate pOH: pOH = -log10[OH-]
    pOH = -log10[0.2] = 0.7

Hydrogen ion concentration, [H+], can be calculated using the following formula:
[H+] = 10-pH

Hydroxide ion concentration, [OH-], can be calculated using the following formula:
[OH-] = 10-pOH

Example:

Find the [H+] of a nitric acid solution with a pH of 3.0
pH= 3.0
[H+] = 10-pH
[H+] = 10-3.0 = 0.001mol L-1

You can check this answer by using the calculated value [H+] in the equation for pH to make sure you arrive at the original pH
pH = -log10[H+]
pH = -log10[0.001] = 3
We get the same value for pH using the calculated value for [H+], so the calculated value for [H+] is correct.

Example:

Find the [OH-] of a sodium hydroxide solution with a pOH of 1
pOH = 1
[OH-] = 10-pOH
[OH-] = 10-1 = 0.1 mol L-1

You can check this answer by using the calculated value [OH-] in the equation for pOH to make sure you arrive at the original pOH
pOH = -log10[OH-]
pOH = -log10[0.1] = 1
We get the same value for pOH using the calculated value for [OH-], so the calculated value for [OH-] is correct.

pH + pOH = 14

Example A(1):

Find the pH of a solution of sodium hydroxide
that has a pOH of 2

pH = 14 - pOH
pH = 14 - 2 = 12

Example B(1):

Find the pOH of a solution of hydrochloric acid
that has a pH of 3.4

pOH = 14 - pH
pOH = 14 - 3.4 = 10.6

Example A(2):

Find the [H+] in a solution of sodium hydroxide
that has a pOH of 1

  • Calculate the pH
    pH = 14 - pOH
    pH = 14 - 1 = 13
  • Calculate [H+]
    [H+] = 10-pH
    [H+] = 10-13 = 10-13mol L-1

Example B(2):

Find the [OH-] of a sulfuric acid solution
with a pH of 3

  • Calculate the pOH
    pOH = 14 - pH
    pOH = 14 - 3 = 11
  • Calculate [OH-]
    [OH-] = 10-pOH
    [OH-] = 10-11 = 10-11 mol L-1

Example A(3):

Find the pH of 0.2mol L-1 sodium hydroxide

  • Write the equation for the dissociation of NaOH:
    NaOH -----> Na+(aq) + OH-(aq)
  • Use the equation to find [OH-]:
    0.2mol L-1 NaOH produces 0.2mol L-1 OH- since NaOH is a strong base that fully dissociates
  • Calculate the pOH:
    pOH = -log10[OH-]
    pOH = -log10[0.2] = 0.7
  • Calculate pH:
    pH = 14 - pOH
    pH = 14 - 0.7 = 13.3

Example B(3):

Find the pOH of 0.2mol L-1 sulfuric acid

  • Write the equation for the dissociation of H2SO4:
    H2SO4 -----> 2H+(aq) + SO42-(aq)
  • Use the equation to find [H+]:
    0.2 mol L-1 H2SO4 produces 2 x 0.2 = 0.4 mol L-1 H+ since H2SO4 is a strong acid that fully dissociates
  • Calculate the pH:
    pH = -log10[H+]:
    pH = -log10[0.4] = 0.4
  • Calculate pOH:
    pOH = 14 - pH
    pOH = 14 - 0.4 = 13.6

 

 

Thanks

Anurag Kishore